OpenStudy (anonymous):

eleven over twelve equals b over twelve ? A.1,584 B. 0.08 C. 10 D. 11

OpenStudy (anonymous):

11/12=b/12

OpenStudy (unklerhaukus):

b=?

OpenStudy (anonymous):

then 12*11/11=b

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

multiply both sides of the equation by twelve

OpenStudy (anonymous):

answer is 11

OpenStudy (aravindg):

11 :)

OpenStudy (anonymous):

lol thanks everyone :)